consider an Island country called Arcadia assume that the aggregate price level is constant, the interest rate is fixed and there are no taxes in this country. according to scenario 1 assume that an increase of 300 in autonomous investment leads to an increase of 750 in equilibrium income. if the MPI is 1/10, the MPS must be?

A. 0.60

B. 0.50

C.0.30

D.0.25
Answer #1

Answer

Multiplier =1/(MPS+MPI)
and also
Multiplier =change in income/change in autonomous spending
multiplier =750/300
=2.5

then from the first equation

2.5=1/(MPS+0.1)
2.5MPS+0.25=1
2.5MPS=0.75
MPS=0.75/2.5=0.3

the MPS is 0.3

Option C